# IBNorm: Information-Bottleneck Inspired Normalization for Representation Learning

### v1

@@ 4.2 IB-Inspired Normalization
In the previous subsection, we identified a key limitation of existing normalization methods: their variance-centric design does not guarantee that intermediate representations preserve task-relevant information. Guided by the IB principle, our goal is to design a normalization operator that explicitly reshapes activations into information-preserving forms, rather than merely controlling their variance.
To this end, we propose an IB-inspired normalization. Unlike conventional variance-centric approaches, which manipulate only first- and second-order activation statistics via standardizing activations to have zero mean and unit variance, our approach introduces an extra compression operation into the normalization operation (NOP) $\psi$ that acts on higher-order statistics. Specifically, $\psi$ compresses activations toward their mean in a controlled manner, thereby increasing local kurtosis and inducing sparsity in activations across instances. Prior work has shown that mean-centered and sparse representations, where most activations concentrate around the mean, often exhibit stronger generalization since redundant and task-unrelated information are removed (Olshausen & Field, 1997; Ranzato et al., 2007; Bengio et al., 2013; Zhang et al., 2018; Guo et al., 2019). Motivated by this insight, our compression operation suppresses variability in activation tails, reshaping the higher-order distribution of activations rather than merely rescaling and shifting activations. More importantly, we find this design aligns normalization with the IB principle, preserving target-relevant information $I(Y;T_{l})$ while suppressing irrelevant task-nuisance information $I(T_{l-1};T_{l})$.
@@
Here we introduce the normalization operation (NOP) $\psi$ in our IBNorm. Let $X$ be hidden activations from a given layer, and let $\bm{x}=\{x_{i}\}_{i=1}^{H}$ represent a sample of $H$-dimensional activations. We define the compression operation $s_{\lambda}(\cdot;\lambda):\mathbb{R}^{d}\to\mathbb{R}^{d}$ as
@@
The property (7) reduces the spread of the activation distribution and increases local kurtosis and sparsity in activations by pushing more activations toward the mean. This sparsity effectively suppresses variability in the tail regions, which often encode task-irrelevant fluctuations in the high-dimensional activations (Olshausen & Field, 1997; Hyvärinen & Oja, 2000). As a result, the compression operation decreases task-nuisance information with the input, $I(T_{l-1};T_{l})$, while preserving the main mass of the distribution that carries task-relevant information, $I(Y;T_{l})$. In other words, by promoting sparse and mean-centered activations, compression encourages representations that retain task relevant information for the downstream task while attenuating task-nuisance information, consistent with the IB principle, which is also empirically validated in Sec. 4.3.
We instantiate three representative forms of the compression function $f_{\lambda}$: IBNorm-S, IBNorm-L, and IBNorm-T. These functionals yield different compression behaviors:
$f_{\lambda}(|x_{i}-\mu|)$ $={|x_{i}-\mu|}/{\lambda},$ (IBNorm-S, linear compression) (8)
$f_{\lambda}(|x_{i}-\mu|)$ $=\ln\!\left(1+{|x_{i}-\mu|}/{\lambda}\right),$ (IBNorm-L, logarithmic compression) (9)
$f_{\lambda}(|x_{i}-\mu|)$ $=\tanh\!\left({|x_{i}-\mu|}/{\lambda}\right).$ (IBNorm-T, hyperbolic tangent compression) (10)
For these functionals, the compression ratio $\alpha_{\lambda}$ in Eqn. (7) becomes $\alpha_{\lambda}=1/\lambda$, guaranteeing the bounded compression property whenever $\lambda\geq 1$. So $\lambda$ controls the compression strength: larger $\lambda$ enforces stronger compression toward the mean $\mu$, whereas smaller $\lambda$ lead to broader activations. By suppressing the tails of the activation distribution, this reduces task-nuisance information while preserving task-relevant information. See more rationale about compression operation in Appendix C.3.
@@
Building on the foundation of LayerNorm, we introduce IBNorm, which integrates an explicit compression operation into the normalization pipeline. This design preserves the well-known benefits of LayerNorm—training stability and acceleration and architectural compatibility—while also incorporating the information-theoretic advantages of the IB principle. Finally, we derive three varaints: IBNorm-S, IBNorm-L, and IBNorm-T.
@@
This hybrid procedure, summarized in Algorithm 1 of Appendix A, yields a normalization strategy explicitly guided by the IB principle. By construction, IBNorm goes beyond variance normalization: it acts as an information filter, enhancing sufficiency by retaining predictive information while improving minimality through compression of irrelevant information in the input. In this way, IBNorm bridges the gap between practical optimization benefits and information-theoretic optimality.
4.3 Theoretical Justification of IBNorm
@@
Information-theoretic superiority of IBNorm. To analyze the impact of normalization, we consider a simplified setting with two one-layer neural networks, $f_{\text{S}}=\Phi_{\text{s}}\circ h$ and $f_{\text{IB}}=\Phi_{\text{IB}}\circ h$, that share the same feature extractor $h$ but differ in their normalization layers: $\Phi_{s}\equiv\eta\circ\psi\circ\zeta$ in the baseline network $f_{\text{s}}$ and $\Phi_{IB}\equiv\eta\circ\psi\circ s_{\lambda}\circ\zeta$ in the network $f_{\text{IB}}$ with IBNorm.
@@
Given the sample dataset $S=\{x_{i},y_{i}\}_{i=1}^{M}\sim\mathbb{P}(X,Y)$ containing $M$ samples, we define the empirical IB value as $\widehat{\text{IB}}_{\text{S}}(T_{1}):=\hat{I}_{\text{S}}(Y;T_{1})-\beta\hat{I}_{\text{S}}(X;T_{1})$ based on Eqn. (5). Because $\eta$ is linear and invertible, it does not affect the mutual information terms; therefore, it suffices to analyze the IB value of intermediate features $\tilde{T}_{s}:=\psi\circ\zeta\circ h(X)$ and $\tilde{T}_{IB}:=\psi\circ s_{\lambda}\circ\zeta\circ h(X)$. Based on the bounded compression property of IBNorm, we can establish the following relationship.
Theorem 1.
[IB Value] For any hyperparameter $\beta\in[0,1]$ and the sample dataset $S\sim\mathbb{P}(X,Y)$ of size $M$, we have
[…]
@@
The proof is provided in Appendix E.1. Theorem 1 establishes the theoretical advantage of IBNorm: its compressed features yield higher IB values than those from normalization with standardization, such as LN and BN. Intuitively, IBNorm better preserves label-relevant information $I(Y;\tilde{T}_{IB})$ while suppressing task-nuisance information $I(X;\tilde{T}_{IB})$, leading to more efficient and predictive representations. Crucially, this benefit holds uniformly across all trade-off parameters $\beta>0$, underscoring the robustness of IBNorm in balancing compression and informativeness.
@@
This result aligns with our empirical observation. We train Llama-130M (Touvron et al., 2023b) and GPT-2 small (Radford et al., 2019) with frozen parameters except the normalization layers on C4 (Raffel et al., 2020) and OpenWebText (Gokaslan & Cohen, 2019), respectively. As shown in Fig. 2(a), IBNorm consistently achieves higher token-level IB values ($\beta=1$) compared to variance-centric normalizations, confirming that the compression operation directly improves information-theoretic optimality. See details in token-level IB value calculation in Appendix G.
@@
Generalization superiority of IBNorm. We next extend the analysis to deep multi-layer networks. Based on the one-layer case, we expect IBNorm to increase IB values layer by layer, thereby enhancing both expressiveness and compression throughout the hierarchy. To formalize this, we study the generalization gap which measures difference between the expected and empirical-training loss:
$\text{gen}(S):=\mathbb{E}_{X,Y}[\mathcal{L}(f(X),Y)]-\tfrac{1}{M}\sum\nolimits_{i=1}^{M}\mathcal{L}(f(x_{i}),y_{i}),$ (12)
@@
where $\mathcal{L}$ denotes a per-sample loss. As shown in prior works (Zhang et al., 2016; Neyshabur et al., 2017; Hoffer et al., 2017; Jakubovitz et al., 2019), a smaller $\text{gen}(S)$ indicates improved generalization performance and robustness both empirically and theoretically.
@@
With probability at least $1-\delta$ over training set $S$ of size $M$, the generalization error of a network $f_{\text{o}}$ satisfies
$\text{gen}(S;f_{\text{o}})\;\leq\;U_{\text{o}}:=\sum\nolimits_{l=1}^{L}U_{l}^{\text{o}},\ \ \text{with}\ \ U_{l}=\sqrt{\tfrac{I(X;\tilde{T}_{l}^{\text{o}}|Y)+C(S,f_{\text{o}})+\log(\frac{2}{\delta})}{M}},$
where $\tilde{T}^{o}_{l}$ is the intermediate representation at layer $l$ before normalization representation recovery $\eta$ and $C(S,f_{\text{o}})$ is a term depending on the network $f_{\text{o}}$ and the training set $S$. Here $f_{\text{o}}$ can be the network $f_{\text{S}}$ using standard normalization like LN and BN, or our network $f_{\text{IB}}$ with our IBNorm.
See Appendix E.5 for the proof. Corollary 2 establishes that the overall generalization gap depends on the cumulative conditional mutual information across layers. This quantity,
$I(X;\tilde{T}_{l}^{\text{o}}|Y)\;=\;I(X;\tilde{T}_{l}^{\text{o}})-I(Y;\tilde{T}_{l}^{\text{o}})\;=\;-\text{IB}(\tilde{T}_{l}^{\text{o}})\quad(\beta=1),$
measures the amount of task-irrelevant input information preserved in the representations.
By Theorem 1, IBNorm increases the IB value of the population distribution at each layer, thereby reducing $I(X;\tilde{T}_{l}^{\text{IB}}|Y)$ compared to $I(X;\tilde{T}_{l}^{\text{S}}|Y)$. Summing across layers, we obtain $U_{\text{IB}}\;\leq\;U_{\text{S}},$ which shows that the generalization bound of the IBNorm network is tighter than that of the standard network. This implies that $f_{\text{IB}}$ achieves superior generalization performance compared to $f_{\text{S}}$. In Fig. 2(b), we follow the setting in Fig. 2(a) to train Llama-130M and GPT-2 small. When evaluated by LLM Leaderboard II (Fourrier et al., 2024), using IBNorm can improve generation performance on downstream tasks, consistent with the information-theoretic analysis (Kawaguchi et al., 2023).
@@
Leaderboard II Leaderboard I
Model Normalization IFEval BBH MATH GPQA MUSR MMLU-PRO AVG AVG
Llama 60M
LayerNorm 0.1479 0.2970 0.0000 0.2576 0.3426 0.1071 0.1920 0.2912
@@
5 Experiment
We evaluate three variants—IBNorm-S, IBNorm-L, and IBNorm-T—across LLM pretraining and vision classification, comparing against LN (Ba et al., 2016), BN (Ioffe & Szegedy, 2015), RMSNorm (Zhang & Sennrich, 2019), and NormalNorm (Eftekhari & Papyan, 2025).
LLM pretraining. We pretrain the LLaMA series (Touvron et al., 2023b) (60M-1B) on C4 (Raffel et al., 2020), following (Lialin et al., 2023; Zhao et al., 2024), and GPT-2 Small (124M) and Medium (355M) on OpenWebText (Gokaslan & Cohen, 2019), following the Sophia setup (Liu et al., 2023) with the nanoGPT implementation (Karpathy, 2022). See details in Appendix F.1. We report performance on LLM Leaderboards I & II (Beeching et al., 2023; Fourrier et al., 2024).
Vision training. We train ResNet-18 on CIFAR-10 (Krizhevsky et al., 2009) and ResNet-50 and ViT on ImageNet (Deng et al., 2009), following the setup (Eftekhari & Papyan, 2025). Both top-1 and top-5 accuracies are reported. Detailed settings are provided in Appendix F.2. Extended results of ViT-S/16 and ViT-B/16 (Dosovitskiy et al., 2020) on ImageNet can be found in Appendix C.2.
5.1 Main Results
@@
Leaderboard II Leaderboard I
Model Normalization IFEval BBH MATH GPQA MUSR MMLU-PRO AVG AVG
GPT-2 124M
LayerNorm 0.1010 0.2809 0.0000 0.2718 0.3624 0.1158 0.1886 0.2887
@@
Table 3: Accuracy (%) on image classification tasks. We report BatchNorm’s results for CNNs (ResNet18/50) since it significantly outperforms LayerNorm on CNNs, and present LayerNorm’s results on Transformers (ViT) because it performs much better than BatchNorm.
ResNet18 on CIFAR-10 ResNet50 on ImageNet ViT on ImageNet
Normalization Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
@@
Results on LLM. Considering the space limitation and the fact that LLM Leaderboard II benchmark can more effectively evaluate the performance and is more widely used now, we defer the results of each task in Leadboard I benchmark to Appendix C.1, and only report its task average result.
As shown in Tab. 1, IBNorm consistently outperforms variance-centric normalization methods such as LN and RMSNorm. On LLaMA-60M, IBNorm-L achieves average scores of 0.2955 (Leaderboard I) and 0.2055 (Leaderboard II), improving over LN by 1.47% and 7.03%, and over RMSNorm by 2.11% and 6.81%. On LLaMA-350M, IBNorm-L reaches 0.3101 (Leaderboard I) and 0.2140 (Leaderboard II), i.e. 1.27% and 6.84% over LN, and 2.03% and 9.51% over RMSNorm. On LLaMA-1B, IBNorm-T achieves the highest scores: 0.3199 on Leaderboard I and 0.2162 on Leaderboard II. Similarly, Tab. 2 also shows that on GPT-2 (355M), IBNorm-L attains 0.3035 (Leaderboard I) and 0.2124 (Leaderboard II), yielding improvements of 2.88% and 9.95% over RMSNorm.
These results demonstrate that IBNorm enhances representation quality in large language models, yielding more predictive and generalizable representations. Further analysis of mutual information and token-level IB values in Appendix C.4 confirms that IBNorm promotes representations that retain task-relevant information while suppressing nuisances, explaining the observed empirical gains.
Results for vision classification. Tab. 3 shows that IBNorm also improves performance across vision models and datasets. On CIFAR-10, ResNet18 with IBNorm-L achieves 95.57% top-1 accuracy, surpassing BN and NormalNorm by 1.92% and 1.56%. On ImageNet, ResNet50 with IBNorm-L improves over BN and NormalNorm by 3.03% and 1.89%. For ViT, IBNorm-L yields top-1 accuracy gains of 5.29%, 5.80%, and 1.58% over LN, RMSNorm, and NormalNorm, respectively.

### v2

@@ 4.2 IB-Inspired Normalization
In Sec. 4.1, we identified a key limitation of existing normalization methods: their variance-centric design does not guarantee that intermediate representations preserve task-relevant information. Guided by the IB principle, our goal is to design a normalization operator that explicitly reshapes activations into information-preserving forms, rather than merely controlling their variance.
To this end, we propose an IB-inspired normalization. Unlike conventional variance-centric approaches, which manipulate only first- and second-order activation statistics via standardizing activations to have zero mean and unit variance, our approach introduces an extra compression operation into the normalization operation (NOP) $\psi$ that acts on higher-order statistics. Specifically, $\psi$ compresses activations toward their mean in a controlled manner, thereby increasing local kurtosis and inducing sparsity in activations across instances. Prior work has shown that mean-centered and sparse representations, where most activations concentrate around the mean, often exhibit stronger generalization by filtering redundant and task-nuisance variability from the tail regions (Olshausen and Field, 1997; Ranzato et al., 2007; Bengio et al., 2013; Zhang et al., 2018; Guo et al., 2019). Motivated by this insight, our compression operation suppresses variability in activation tails, reshaping the higher-order distribution of activations rather than merely rescaling and shifting activations. Importantly, we find this design aligns normalization with the IB principle, preserving target-relevant information $I(Y;T_{l})$ while suppressing irrelevant task-nuisance information $I(T_{l-1};T_{l})$ (Sec. 4.3).
@@
Here we introduce the normalization operation (NOP) $\psi$ in our IBNorm. Let $X$ be hidden activations from a given layer, and let $\bm{x}=\{x_{i}\}_{i=1}^{H}$ represent a sample of $H$-dimensional activations. We define the compression operation $s_{\lambda}(\cdot;\lambda):\mathbb{R}^{d}\to\mathbb{R}^{d}$:
@@
The property (7) increases local kurtosis and reduces the spread of the distribution in activations by compressing activations toward the mean. This compression suppresses variability in the tail regions, which often encode task-irrelevant fluctuations in the high-dimensional activations (Olshausen and Field, 1997; Hyvärinen and Oja, 2000). As a result, the compression operation decreases task-nuisance information with the input, $I(T_{l-1};T_{l})$, while preserving the main mass of the distribution that carries task-relevant information, $I(Y;T_{l})$. In other words, by promoting mean-centered activations and inducing effective sparsification in IBNorm, compression encourages representations that retain task-relevant information for the downstream task while attenuating task-nuisance information, consistent with the IB principle, which is also empirically validated in Sec. 4.3.
While the bounded compression property (Eq. 7) establishes the fundamental condition for $f_{\lambda}$, the rate of tail decay remains an important degree of freedom in its design (Chen et al., 2025; Ash and Doléans-Dade, 2000). To systematically explore the design space, we instantiate three representative functionals—IBNorm-S, IBNorm-L, and IBNorm-T—that span a spectrum of tail suppression strength. As visualized in Figures 46, these variants apply increasingly stronger suppression to the activation tails: IBNorm-S is the mildest, IBNorm-L is intermediate, and IBNorm-T is the most aggressive. These structural differences manifest in the induced entropy of the representation distribution, leading to different empirical performance. See Appendix C.6 for detailed analysis. Their specific formulations are:
$f_{\lambda}(|x_{i}-\mu|)$ $={|x_{i}-\mu|}/{\lambda},$ (IBNorm-S) (8)
$f_{\lambda}(|x_{i}-\mu|)$ $=\ln\!\left(1+{|x_{i}-\mu|}/{\lambda}\right),$ (IBNorm-L)
$f_{\lambda}(|x_{i}-\mu|)$ $=\tanh\!\left({|x_{i}-\mu|}/{\lambda}\right).$ (IBNorm-T)
For these functionals, the compression ratio $\alpha_{\lambda}$ in Eqn. (7) becomes $\alpha_{\lambda}=1/\lambda$, guaranteeing the bounded compression property whenever $\lambda\geq 1$. Thus, $\lambda$ controls the compression strength: larger $\lambda$ enforces stronger compression toward the mean $\mu$, whereas smaller $\lambda$ lead to broader activations. This mechanism acts as a structural prior to reduce task-nuisance information while preserving task-relevant information.
@@
Building on the foundation of LayerNorm, we introduce IBNorm, which integrates an explicit compression operation into the normalization pipeline. This design preserves the well-known benefits of LayerNorm—training stability and acceleration and architectural compatibility—while also incorporating the information-theoretic advantages of the IB principle. Finally, we derive three variants: IBNorm-S, IBNorm-L, and IBNorm-T.
Formally, given an activation input $\bm{x}$, IBNorm applies the following sequence of operations: first, the NAP operator $\zeta$ partitions features in the same way as LayerNorm; next, a compression operator $s_{\lambda}$ reduces nuisance variability; then, the NOP operator $\psi$ standardizes activations and induces effective sparsity in representation based on compression;
@@
This hybrid procedure, summarized in Algorithm 1 of Appendix A.3, yields a normalization strategy explicitly guided by the IB principle. We highlight that this differs from the function of nonlinear activations (see Appendix A.1). By construction, IBNorm goes beyond variance normalization: it acts as an information filter, enhancing sufficiency by retaining predictive information while improving minimality through compression of irrelevant information in the input. In this way, IBNorm bridges the gap between practical optimization benefits and information-theoretic optimality.
@@ 4.3 Theoretical Justification of IBNorm
Information-theoretic superiority of IBNorm. To analyze normalization, we consider a simplified setting with two one-layer neural networks, $f_{\text{S}}=\Phi_{\text{s}}\circ h$ and $f_{\text{IB}}=\Phi_{\text{IB}}\circ h$, that share the same feature extractor $h$ but differ in their normalization layers: $\Phi_{s}\equiv\eta\circ\psi\circ\zeta$ in the baseline network $f_{\text{S}}$ and $\Phi_{IB}\equiv\eta\circ\psi\circ s_{\lambda}\circ\zeta$ in the network $f_{\text{IB}}$ with IBNorm.
@@
Given the sample dataset $S=\{x_{i},y_{i}\}_{i=1}^{M}\sim\mathbb{P}(X,Y)$ containing $M(<\infty)$ samples, we define the empirical IB value as $\widehat{\text{IB}}_{\text{S}}(T_{1}):=\hat{I}_{\text{S}}(Y;T_{1})-\beta\hat{I}_{\text{S}}(X;T_{1})$ based on Eqn. (5); therefore, we aim to analyze the empirical IB value of features $T_{s}:=f_{\text{S}}(X)$ and $T_{IB}:=f_{\text{IB}}(X)$.
Based on the bounded compression property of IBNorm, we can establish the following relationship.
Theorem 1 (IB Value).
For any hyperparameter $\beta\in[0,1]$ and the sample dataset $S\sim\mathbb{P}(X,Y)$ of size $M$, we have
$\widehat{\text{IB}}_{S}(T_{IB})\geq\widehat{\text{IB}}_{S}(T_{s})\qquad\text{almost surely}.$ (10)
@@
The proof is provided in Appendix E.1. Theorem 1 establishes the theoretical advantage of IBNorm: its compressed features yield higher IB values than those from normalization with standardization, such as LN and BN. Intuitively, IBNorm better preserves label-relevant information $I(Y;T_{IB})$ while suppressing task-nuisance information $I(X;T_{IB})$, leading to more efficient and predictive representations. Crucially, this benefit holds uniformly across all trade-off parameters $\beta>0$, underscoring the robustness of IBNorm in balancing compression and informativeness.
@@
This result aligns with our empirical results. We train Llama-130M and GPT-2 small with frozen parameters except the normalization layers on C4 (Raffel et al., 2020) and OpenWebText (Gokaslan and Cohen, 2019), respectively. As shown in Fig. 2(a), IBNorm consistently achieves higher token-level IB values ($\beta=1$) compared to variance-centric normalizations, confirming that the compression operation improves information-theoretic optimality. See details in token-level IB value calculation in Appendix G.
@@
Generalization superiority of IBNorm. We next extend the analysis to deep multi-layer networks. Motivated by the one-layer result, we expect that IBNorm can increase IB values of multi-layer networks (Appendix C.7), thereby enhancing both expressiveness and compression. To formalize this, we study the generalization gap which measures difference between the expected and empirical-training loss:
$\text{gen}(S)=\mathbb{E}_{X,Y}[\mathcal{L}(f(X),Y)]-\tfrac{1}{M}\sum\nolimits_{i=1}^{M}\mathcal{L}(f(x_{i}),y_{i}),$ (11)
@@
where $\mathcal{L}$ denotes a per-sample loss. As shown in prior works (Zhang et al., 2016; Neyshabur et al., 2017; Hoffer et al., 2017; Jakubovitz et al., 2019), a smaller $\text{gen}(S)$ indicates improved generalization performance and robustness theoretically. Without loss of generality, we assume $\beta=1$.
@@
With probability at least $1-\delta$ over training set $S$ of size $M<\infty$, the generalization error of a L-layer network $f_{\text{o}}$ satisfies
$\text{gen}(S;f_{\text{o}})\;\leq\;U_{\text{o}}=\sum\nolimits_{l=1}^{L}U_{l}^{\text{o}},$ (12)
$\text{with}\ \ U^{\text{o}}_{l}=\sqrt{\tfrac{-\text{IB}(T_{l}^{\text{o}})+C_{S}+\log(\frac{2}{\delta})}{M}},$
where $T^{\text{o}}_{l}$ is the intermediate representation at layer $l$, and $C_{S}$ is a term depending on the cardinality of the training set $S$. Here $f_{\text{o}}$ can be the network $f_{\text{S}}$ using standard normalization like LN and BN, or the network $f_{\text{IB}}$ with IBNorm.
See Appendix E.5 for the proof. Corollary 2 establishes that the overall generalization gap depends on the quantity $\text{IB}(T_{l}^{\text{o}})$ which measures the amount of task-irrelevant input information preserved in the representations.
By Theorem 1, IBNorm increases the IB value of the population distribution at each layer, thereby reducing $\text{IB}(T_{l}^{\text{IB}})$ compared to $\text{IB}(T_{l}^{\text{S}})$. Summing across layers, we obtain $U_{\text{IB}}\;\leq\;U_{\text{S}},$ which shows that the generalization bound of the IBNorm network is tighter than that of the standard network. This implies that $f_{\text{IB}}$ achieves superior generalization performance compared to $f_{\text{S}}$. In Fig. 2(b), we follow the setting in Fig. 2(a) to train Llama-130M and GPT-2 small. Evaluation under LLM Leaderboard II shows that IBNorm improves generalization performance, consistent with the information-theoretic  (Kawaguchi et al., 2023) and empirical studies (Zhu et al., 2025; Chen et al., 2025).
5 Experiment
We evaluate three variants—IBNorm-S, IBNorm-L, and IBNorm-T—across LLM pretraining and vision classification, comparing against LN (Ba et al., 2016), BN (Ioffe and Szegedy, 2015), RMSNorm (Zhang and Sennrich, 2019), and NormalNorm (Eftekhari and Papyan, 2025).
LLM pretraining. We pretrain LLaMA series (Touvron et al., 2023b) (60M-1B) on C4 (Raffel et al., 2020), following (Lialin et al., 2023; Zhao et al., 2024), and GPT-2 (124M and 355M) on OpenWebText (Gokaslan and Cohen, 2019), following the Sophia setup (Liu et al., 2023) with the nanoGPT implementation (Karpathy, 2022). See details in Appendix F.1. We report performance on LLM Leaderboards I & II (Beeching et al., 2023; Fourrier et al., 2024).
Vision training. We train ResNet-18 on CIFAR-10 (Krizhevsky et al., 2009) and ResNet-50 and ViT on ImageNet (Deng et al., 2009), following the setup (Eftekhari and Papyan, 2025). Both top-1 and top-5 accuracies are reported. Detailed settings are provided in Appendix F.2. Extended results of ViT-S/16 and ViT-B/16 (Dosovitskiy et al., 2020) on ImageNet can be found in Appendix C.5.
5.1 Main Results
@@
Model Normalization Leaderboard II Leaderboard I
IFEval ($\uparrow$) BBH ($\uparrow$) MATH ($\uparrow$) GPQA ($\uparrow$) MUSR ($\uparrow$) MMLU-PRO ($\uparrow$) AVG ($\uparrow$) AVG ($\uparrow$)
Llama 60M LayerNorm 0.1479 0.2970 0.0000 0.2576 0.3426 0.1071 0.1920 0.2912
@@
Model Normalization Leaderboard II Leaderboard I
IFEval ($\uparrow$) BBH ($\uparrow$) MATH ($\uparrow$) GPQA ($\uparrow$) MUSR ($\uparrow$) MMLU-PRO ($\uparrow$) AVG ($\uparrow$) AVG ($\uparrow$)
GPT-2 124M LayerNorm 0.1010 0.2809 0.0000 0.2718 0.3624 0.1158 0.1886 0.2887
@@
Table 3: Accuracy (%) on image classification tasks. We report BatchNorm’s results for CNNs (ResNet18/50) since it outperforms LayerNorm on CNNs, and present LayerNorm’s results on Transformers (ViT) because it performs much better than BatchNorm.
Normalization ResNet18 on CIFAR-10 ResNet50 on ImageNet ViT on ImageNet
Top-1 Top-5 Top-1 Top-5 Top-1 Top-5
@@
Results on LLM. Due to the space limitation, we defer the results of each task in Leaderboard I to Appendix C.1 and report results with statistical significance in Appendix C.2.
As shown in Tab. 1, IBNorm outperforms variance-centric normalization methods such as LN and RMSNorm. On LLaMA-60M, IBNorm-L achieves average scores of 0.2955 (Leaderboard I) and 0.2055 (Leaderboard II), improving over LN by 1.47% and 7.03%, and over RMSNorm by 2.11% and 6.81%. On LLaMA-350M, IBNorm-T reaches 0.3101 (Leaderboard I) and 0.2140 (Leaderboard II), i.e. 1.27% and 6.84% over LN, and 2.51% and 6.46% over RMSNorm. On LLaMA-1B, IBNorm-T achieves the highest scores: 0.3199 on Leaderboard I and 0.2162 on Leaderboard II. Similarly, Tab. 2 also shows that on GPT-2 (355M), IBNorm-L attains 0.3035 (Leaderboard I) and 0.2124 (Leaderboard II), yielding improvements of 2.88% and 9.95% over RMSNorm.
These results demonstrate that IBNorm enhances representation quality in large language models, yielding more predictive and generalizable representations. Further analysis of mutual information and token-level IB values in Appendix C.7 confirms that IBNorm promotes representations that retain task-relevant information while suppressing nuisances, explaining the observed empirical gains. However, IBNorm shows some failure cases (see Appendix C.1).
Results for vision classification. Tab. 3 shows that IBNorm also improves performance across vision models and datase-
Table 4: Ablation study of Llama 60M with normalization using different $\lambda$ evaluated on LLM Leaderboard I and LLM Leaderboard II. We defer the detailed results of each task to Appendix C.4.
Normalization $\lambda$ Leaderboard I Leaderboard II
AVG ($\uparrow$) AVG ($\uparrow$)
IBNorm-S 0.5 0.2898 0.1977
4 0.2952 0.1987
8 0.2918 0.1981
IBNorm-L 0.5 0.2863 0.2018
4 0.2955 0.2055
8 0.2921 0.1881
IBNorm-T 0.5 0.2881 0.1984
4 0.2937 0.2025
8 0.2898 0.2005
Table 5: Ablation study of Llama 60M evaluated on LLM Leaderboard I and LLM Leaderboard II. Due to space limitation, we defer the detailed results of each task to Appendix C.3.
Normalization Leaderboard I Leaderboard II
AVG ($\uparrow$) AVG ($\uparrow$)
IBNorm-S* 0.2952 0.1978
IBNorm-S** 0.2881 0.1945
IBNorm-T* 0.2936 0.1956
IBNorm-T** 0.2831 0.1940
IBNorm-L* 0.2905 0.1967
IBNorm-L** 0.2859 0.1934
ts. On CIFAR-10, ResNet18 with IBNorm-L achieves 95.57% top-1 accuracy, surpassing BN and NormalNorm by 1.92% and 1.56%. On ImageNet, ResNet50 with IBNorm-L improves over BN and NormalNorm by 3.03% and 1.89%. For ViT, IBNorm-L yields top-1 accuracy gains of 5.29%, 5.80%, and 1.58% over LN, RMSNorm, and NormalNorm.